# Does Unlearning Truly Unlearn? A Black Box Evaluation of LLM Unlearning Methods

### v2

@@ Abstract
Large language model unlearning aims to remove harmful information that LLMs have learnt to prevent their use for malicious purposes. LLMU and RMU have been proposed as two methods for LLM unlearning, achieving impressive results on unlearning benchmarks. We study in detail the efficacy of these methods by evaluating their impact on general model capabilities on the WMDP benchmark as well as a biology benchmark we create. Our experiments show that RMU generally leads to better preservation of model capabilities, for similar or better unlearning. We further test the robustness of these methods and find that doing 5-shot prompting or rephrasing the question in simple ways can lead to an over ten-fold increase in accuracy on unlearning benchmarks. Finally, we show that training on unrelated data can almost completely recover pre-unlearning performance, demonstrating that these methods fail at truly unlearning. The code is available at: https://github.com/JaiDoshi/Knowledge-Erasure.
@@ 1 Introduction
Two recent methods for LLM unlearning are LLMU (Yao et al., 2023) and RMU (Li et al., 2024a). LLMU performs gradient ascent on the harmful data by taking the negative of the loss function. RMU trains particular layers of the model towards a random control vector on harmful data. Both of these methods employ an additional “retain set" which consists of benign data used to preserve normal capabilities.
Although these works show that the methods are effective at unlearning, they are limited in the scope of their evaluation. We extend the evaluation to a new biology-focused dataset, use three main metrics to study the effect on general model capabilities, and design robustness tests based on prompting strategies that an adversary may use. These include doing 5-shot prompting instead of zero-shot prompting, and rephrasing the question in different ways, such as in the form of a poem. Applying these simple prompting strategies leads to an increase in accuracy of up to 1750% on unlearning benchmarks, suggesting that the information has not been actually removed from the model. We test this hypothesis by checking the impact of retraining on benign data, and find that this undoes the effect of unlearning and restores harmful capabilities. Figure 1 shows an overview of our approach.
Figure 1: Applying unlearning techniques on the pretrained model makes it give nonsensical outputs to harmful queries. Using robustness tests (in this case adding filler text before the question) or fine-tuning on benign data causes unlearned capabilities to resurface.
@@ 2 Related Work
Liu et al. (2024) expound unlearning effectiveness and utility preservation (along with efficiency) as the main criteria for evaluating unlearning methods. Łucki et al. (2024) focus on the robustness of unlearning, and use methods such as Logit Lens and a modified version of GCG to recover unlearned performance. They show that fine-tuning on just a few unrelated examples is able to almost completely undo the effects of unlearning. The methods they use assume access to the model weights, and so cannot be applied to black box LLMs except for cases where model fine-tuning is possible through API access. Lynch et al. (2024) define eight metrics that can be used to evaluate the robustness of LLM unlearning methods, including translating the queries to another language. We adapt some of these metrics in our work, as well as introduce new ones that can be applied to black box models.
@@ 3.1 Unlearning Methods
Yao et al. (2023) define three losses: the gradient ascent loss $\mathcal{L}_{\mathrm{fgt}}$, $\mathcal{L}_{\mathrm{rdn}}$ which trains the LLM towards a random output for an input that is to be unlearned, and $\mathcal{L}_{\mathrm{nor}}$ which computes the KL divergence between the unlearning model and normal model on the retain set to preserve normal utility. These losses are weighted by hyperparameters $\epsilon_{1}$, $\epsilon_{2}$ and $\epsilon_{3}$ respectively. $\mathcal{L}_{\mathrm{rdn}}$ is better suited for QA datasets, so we use only $\mathcal{L}_{\mathrm{fgt}}$ and $\mathcal{L}_{\mathrm{nor}}$.
@@ 3.2 Datasets
We consider the example task of unlearning all biology information from the model. We believe this task serves as a good proxy to real world applications where a particular domain needs to be forgotten. We scrape all Wikipedia articles under biology related categories, recursively processing all the subcategories within these categories until a certain depth. Although some of the articles may not be relevant, such as biographies of biologists, we keep these articles as a measure to whether the unlearning methods are robust to unclean data. We similarly process randomly chosen non-biology categories to create the retain set. The lists of categories used and the depths of recursive processing are given in Appendix A.2.
@@
We additionally unlearn on the WMDP dataset (Li et al., 2024a), focusing on unlearning the WMDP-Bio and WMDP-Chem subsets, and use Wikitext as the retain set.
@@ 1. Non-generation of harmful responses
We use accuracy on multiple choice benchmarks as the primary metric. For biology unlearning we consider the macro-average accuracy on MMLU (Hendrycks et al., 2021) for biology-related subjects (refer to Appendix A.3 for the list of subjects). For WMDP, we use the provided QA dataset, and evaluate on the Biology and Cyber subsets separately. We use a zero-shot template akin to Li et al. (2024a), but remove the name of the subject as it increases the model’s tendency to refuse to answer. An example question is provided in Appendix A.4. To measure the robustness of unlearning, we use the following two additional tests:
@@
In order to encourage the model to answer the question, we include five examples of answered multiple choice questions of an MMLU subject before the actual question. Appendix A.5.1 contains the list of MMLU subjects used in our experiments and Appendix A.5.2 contains an example question.
@@ 3. Prevention of recovery of unlearned information
Lynch et al. (2024) list fine-tuning on harmful data as one of the metrics to evaluate unlearning. However, this requires the malicious user to have access to a dataset of harmful information. Similar to Łucki et al. (2024), we experiment with using the benign Openwebtext dataset to see if unlearned performance can be recovered by fine-tuning on generic web data.
@@ 5.1 Unlearning and Robustness Tests
Zephyr LLMU 0.074 0.472 0.259 (+247%) 0.535 5-shot High School Physics
Zephyr RMU 0.097 0.386 0.274 (+181%) 0.353 5-shot Professional Law
@@
Llama RMU 0.136 0.342 0.376 (+177%) 0.424 5-shot Jurispudence
@@
Zephyr LLMU 0.185 0.689 0.220 (+19.18%) 0.605 Technical Terms Removed 1
Zephyr RMU 0.146 0.391 0.274 (+87.7%) 0.350 5-shot Professional Law
@@
Llama RMU 0.060 0.336 0 < 35% -
(c) WMDP Cyber dataset
Table 1: Accuracies on the respective multiple choice benchmarks along with the accuracies when the model actually answers the question, as well as the maximum accuracy (except for cases where the accuracy when answered is less than 35%) attained from applying robustness tests and the robustness test responsible for the maximum accuracy.
Table 1 contains the accuracies on the benchmarks, as well as the maximum accuracy attained from applying all the robustness tests. The unlearning methods appear to perform well in all cases, with performance on the benchmarks reduced to under 20% (from up to 71%) in all cases. We observe that the accuracy when the model answers the question is significantly higher for LLMU than RMU.
For the robustness tests, since a model that predicts random answers has an accuracy of ~25%, we also take into consideration the accuracy of the model when it answers the question, i.e. when it outputs an option rather than a random character. When this is less than 35% post applying the tests, we consider the model outputs too close to random chance for them to be useful, and say that unlearning has been successful for the benchmark. We observe that for the Biology dataset, 5-shot prompting is enough to increase accuracy close to or more than 3 times. For WMDP, the unlearned models tend to be more robust, however, there is an increase of over 1100% in accuracy for Llama LLMU on translating to certain lower-resource languages such as Farsi.
@@ 5.2 Preservation of Model Capabilities
We compare the accuracy from the unlearning benchmarks with performance on MMLU and MT-Bench. Plots are contained in Figure 3. In almost all cases, apart from Zephyr RMU on WMDP, there is a noticeable drop in both MMLU and MT-Bench performance post unlearning (>5%). We observe that for the Biology dataset, both methods perform similarly, with the exception that Llama LLMU has an accuracy on MMLU Other of 60.7% as compared to 54.6% for Llama RMU (accuracy on the original model is 64.1%). For the WMDP dataset however, there is a significant degradation in normal performance for LLMU. On the Biology subset with Llama for instance, MMLU accuracy drops from 65.1% to 51.2% and MT-Bench score drops from 8.07 to 5.17.
@@ 5.3 Fine-tuning on Benign Data
Table 3: Accuracies on the unlearning tasks post fine-tuning on Openwebtext.
@@ 6 Conclusions
Overall we find RMU tends to perform better than LLMU, particularly in terms of retaining general model capabilities. It has the additional advantage of being less computationally expensive—requiring fewer steps to unlearn and training on fewer layers. LLMU also tends to significantly increase the perplexity of the model. The results from fine-tuning on Openwebtext support the findings of Łucki et al. (2024) and strongly suggest that both RMU and LLMU learn a filter that makes the model refuse to answer harmful queries rather than actually remove information from the model. Our robustness tests show that this filter can be bypassed using simple prompting techniques and motivate the development of unlearning methods that truly remove information from the model.
@@
We believe the following three points are the main limitations of our work: (1) For LLMU, we did not use $\mathcal{L}_{\mathrm{fgt}}$ as our data was not in QA format. Incorporating this loss might have led to improved results. (2) We observed that training both RMU and LLMU is highly stochastic, and the performance varies significantly between checkpoints for even the same set of hyperparameters. Consequently, the results may have been different had we saved checkpoints more frequently. However it is practically infeasible to save the weights more frequently and run all the evaluation tests every time. (3) Although our robustness tests were successful in most cases, we were unable to recover meaningful performance from Llama RMU on WMDP. Further work can look at developing additional tests that are more effective.

### v3

@@ Abstract
Large language model unlearning aims to remove harmful information that LLMs have learnt to prevent their use for malicious purposes. LLMU and RMU have been proposed as two methods for LLM unlearning, achieving impressive results on unlearning benchmarks. We study in detail the impact of unlearning on LLM performance metrics using the WMDP dataset as well as a new biology dataset we create. We show that unlearning has a notable impact on general model capabilities, with the performance degradation being more significant in general for LLMU. We further test the robustness of the two methods and find that doing 5-shot prompting or rephrasing the question in simple ways can lead to an over ten-fold increase in accuracy on unlearning benchmarks. Finally, we show that training on unrelated data can almost completely recover pre-unlearning performance, demonstrating that these methods fail at truly unlearning. Our methodology serves as an evaluation framework for LLM unlearning methods. The code is available at: https://github.com/JaiDoshi/Knowledge-Erasure.
@@ 1 Introduction
Two recent methods for LLM unlearning are LLMU (Yao et al., 2023) and RMU (Li et al., 2024a). Although these works show that the methods are effective at unlearning, they are limited in the scope of their evaluation and robustness testing. We introduce a new biology-focused dataset that serves as a measure of the ability to unlearn on noisy data; use three main metrics to study the effect of unlearning on general model capabilities; and design robustness tests based on prompting strategies that an adversary may use. These include doing 5-shot prompting and rephrasing the question in different ways, such as in the form of a poem. Applying these simple prompting strategies leads to an increase in accuracy of up to 1750% on unlearning benchmarks, suggesting that the information has not been actually removed from the model. We test this hypothesis by checking the impact of retraining on benign data, and find that this undoes the effect of unlearning and restores harmful capabilities. Figure 1 highlights our robustness testing approach.
Figure 1: Applying unlearning techniques on the pretrained model makes it give nonsensical outputs to harmful queries. Adversarial prompting (in this case adding filler text before the question) or fine-tuning on benign data causes unlearned capabilities to resurface.
Overall, our work develops a framework for evaluating LLM unlearning with a focus on black box methods for robustness testing. The application of this framework to two recent unlearning methods demonstrates deterioration in general model capabilities and ineffectiveness at truly unlearning.
@@ 2 Related Work
Liu et al. (2024) expound unlearning effectiveness and utility preservation (along with efficiency) as the main criteria for evaluating unlearning methods. Łucki et al. (2024) focus on the robustness of unlearning, and use methods such as Logit Lens and a modified version of GCG to recover unlearned performance. They show that fine-tuning on just a few unrelated examples is able to almost completely undo the effects of unlearning. However, the methods they use assume access to the model weights, and so cannot be applied to black box LLMs except for cases where model fine-tuning is possible through API access. Lynch et al. (2024) define eight metrics that can be used to evaluate the robustness of LLM unlearning methods, including translating the queries to another language. We adapt some of these metrics in our work, as well as introduce new ones that can be applied to black box models.
@@ 3.1 Unlearning Methods
Yao et al. (2023) define three losses: the gradient ascent loss $\mathcal{L}_{\mathrm{fgt}}$, $\mathcal{L}_{\mathrm{rdn}}$ which trains the LLM towards a random output for an input that is to be unlearned, and $\mathcal{L}_{\mathrm{nor}}$ which computes the KL divergence between the unlearning model and normal model on a retain set to preserve normal utility. These losses are weighted by hyperparameters $\epsilon_{1}$, $\epsilon_{2}$ and $\epsilon_{3}$ respectively. $\mathcal{L}_{\mathrm{rdn}}$ is better suited for QA datasets, so we use only $\mathcal{L}_{\mathrm{fgt}}$ and $\mathcal{L}_{\mathrm{nor}}$.
@@ 3.2 Datasets
We consider the example task of unlearning all biology information from the model. We believe this task serves as a good proxy to real world applications where a particular domain needs to be forgotten. We scrape all Wikipedia articles under biology-related categories, recursively processing all the subcategories within these categories until a certain depth. Although some of the articles may not be relevant, such as biographies of biologists, we keep these articles to measure the ability to unlearn on unclean data. We similarly process randomly chosen non-biology categories to create the retain set. The lists of categories used and the depths of recursive processing are given in Appendix A.2.
@@
We additionally unlearn on the WMDP dataset (Li et al., 2024a), focusing on unlearning the WMDP-Bio and WMDP-Cyber subsets, and use Wikitext as the retain set.
@@ 1. Non-generation of harmful responses
We use accuracy on multiple choice benchmarks as the primary metric. For biology unlearning we consider the macro-average accuracy on MMLU (Hendrycks et al., 2021) for biology-related subjects (refer to Appendix A.3 for the list of subjects). For WMDP, we use the provided QA dataset, and evaluate on the Biology and Cyber subsets. We use a zero-shot template akin to Li et al. (2024a), but remove the name of the subject as it increases the model’s tendency to refuse to answer. An example question is provided in Appendix A.4. To measure the robustness of unlearning, we use the following two additional tests:
@@
In order to encourage the model to answer the question, we include five examples of answered multiple choice questions of an MMLU subject before the actual question. Appendix A.5.1 contains the list of MMLU subjects used and Appendix A.5.2 contains an example question.
@@ 3. Prevention of recovery of unlearned information
Lynch et al. (2024) list fine-tuning on harmful data as one of the metrics to evaluate unlearning. However, this requires the malicious user to have access to a dataset of harmful information. Similar to contemporary work by Łucki et al. (2024), we experiment with using the benign Openwebtext dataset to see if unlearned performance can be recovered by fine-tuning on generic web data.
@@ 5.1 Unlearning and Robustness Tests
Zephyr LLMU 0.074 0.472 0.259 (+250%) 0.535 5-shot High School Physics
Zephyr RMU 0.097 0.386 0.298 (+207%) 0.312 Translated to Telugu
@@
Llama RMU 0.136 0.342 0.376 (+176%) 0.424 5-shot Jurispudence
@@
Zephyr LLMU 0.185 0.689 0.254 (+37.3%) 0.333 Translated to Bengali
Zephyr RMU 0.146 0.391 0.293 (+101%) 0.321 Translated to Hindi
@@
Llama RMU 0.060 0.336 0.164 (+173%) 0.274 Technical Terms Removed 2
(c) Accuracies on the WMDP Cyber Benchmark.
Table 1: Accuracies on the respective multiple choice benchmarks along with the accuracies when the model actually answers the question, as well as the maximum accuracy attained from applying robustness tests and the robustness test responsible for the maximum accuracy.
Table 1 contains the accuracies on the benchmarks, as well as the maximum accuracy attained from applying all the robustness tests. The unlearning methods appear to perform well, with performance on the benchmarks reduced to under 20% (from up to 71%) in all cases. We observe that the accuracy when the model answers the question is significantly higher for LLMU than RMU.
For the Biology dataset, the robustness tests increase accuracy close to or more than 3 times. For WMDP, the unlearned models tend to be more robust, however, there is an increase of over 1100% in accuracy for Llama LLMU on translating to certain lower-resource languages such as Farsi. Overall, 5-shot prompting and translations are observed to be the most effective robustness tests. We perform a Z-test based on the accuracy when answered to confirm that the model answers are not random.
@@ 5.2 Preservation of Model Capabilities
We compare the accuracy from the unlearning benchmarks with performance on MMLU and MT-Bench. Plots are contained in Figure 3. In almost all cases, apart from Zephyr RMU on WMDP, there is a noticeable drop in both MMLU and MT-Bench performance post unlearning (>5%). We observe that for the Biology dataset, both methods perform similarly, with the exception that Llama LLMU has an accuracy on MMLU Other of 60.7% as compared to 54.6% for Llama RMU (accuracy on the original model is 64.1%). On the WMDP dataset however, there is a significant degradation in normal performance for LLMU. For instance, on the Biology subset with Llama, MMLU accuracy drops from 65.1% to 51.2% and MT-Bench score drops from 8.07 to 5.17, and RMU is Pareto optimal compared to LLMU for both subsets for Llama.
@@ 5.3 Fine-tuning on Benign Data
Table 3: Accuracies on the unlearning benchmarks post fine-tuning on Openwebtext.
@@ 6 Conclusions
Overall we find RMU tends to perform better than LLMU, particularly in terms of retaining general model capabilities. It has the additional advantage of being less computationally expensive—requiring fewer steps to unlearn and training on fewer layers. LLMU also tends to significantly increase the perplexity of the model. The results from fine-tuning on Openwebtext support the findings of Łucki et al. (2024) and extend them to LLMU, and strongly suggest that unlearning methods learn a filter that makes the model refuse to answer harmful queries rather than actually remove information from the model. Our robustness tests show that this filter can be bypassed using simple prompting techniques and motivate the development of unlearning methods that truly remove information from the model.
@@
We believe the following points to be the main limitations of our work: (1) We chose LLMU and RMU as the unlearning methods to experiment on as these were the LLM-specific methods studied by Li et al. (2024a). Future work can cover other methods, although experiments by Łucki et al. (2024) on NPO (Zhang et al., 2024) suggest a similar failure at unlearning. (2) For LLMU, we did not use $\mathcal{L}_{\mathrm{fgt}}$ as our data was not in QA format. Incorporating this loss might have led to improved results. (3) We observed that training both RMU and LLMU is highly stochastic and that performance varies significantly between checkpoints for even the same set of hyperparameters. Consequently, the results may have been different had we saved checkpoints more frequently. However it is practically infeasible to save the weights more frequently and run all the evaluation tests every time.